Solve for x and y
3x - 4y = 32
x + 3y = -11

Answer by mananth(16946)

3 x -4 y = 32 .............1
1 x + 3 y = -11 .............2
Eliminate y
multiply (1)by 3
Multiply (2) by 4
9 x -12 y = 96
4 x + 12 y = -44
Add the two equations
13 x = 52.00
/ 13
x = 4
plug value of x in (1)
3 x + -4 y = 32
12 + -4 y = 32
-4 y = 32 -12
-4 y = 20
y = -5